OpenStudy (anonymous):

Can anyway help me with this problem? √154x^5y^3

OpenStudy (anonymous):

this?\[\sqrt{154x^5y^3}\]

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

x2ysqrt154xy x2 is x squared

OpenStudy (anonymous):

\[\sqrt{154}\cdot x^2\cdot \sqrt{x}\cdot y\cdot\sqrt {y}\]

OpenStudy (agreene):

\[\large x^{\frac52}y^{\frac32}\sqrt{154}\] iff x,y are real and positive.